# Mamba in Speech: Towards an Alternative to Self-Attention

### v4

@@ Abstract
Transformer and its derivatives have achieved success in diverse tasks across computer vision, natural language processing, and speech processing. To reduce the complexity of computations within the multi-head self-attention mechanism in Transformer, Selective State Space Models (i.e., Mamba) were proposed as an alternative. Mamba exhibited its effectiveness in natural language processing and computer vision tasks, but its superiority has rarely been investigated in speech signal processing. This paper explores solutions for applying Mamba to speech processing using two typical speech processing tasks: speech recognition, which requires semantic and sequential information, and speech enhancement, which focuses primarily on sequential patterns. The experimental results exhibit the superiority of bidirectional Mamba (BiMamba) for speech processing to vanilla Mamba. Moreover, experiments demonstrate the effectiveness of BiMamba as an alternative to the self-attention module in Transformer and its derivates, particularly for the semantic-aware task. The crucial technologies for transferring Mamba to speech are then summarized in ablation studies and the discussion section to offer insights for future research. 111*Equal contribution, Our open accessed repository can be found via https://github.com/Tonyyouyou/Mamba-in-Speech
1 Introduction
@@
Transformer-based models [1] have shone brightly across various domains in machine learning, including computer vision (CV) [2, 3, 4], natural language processing (NLP) [5, 6, 7], and speech processing [8, 9]. This success is linked to the multi-head self-attention (MHSA) module, which facilitates the representation of intricate data structures within a specific context window. However, the self-attention mechanism encounters a challenge with computational complexity, which grows quadratically as the size of the context window increases. In speech tasks, the window typically encompasses an entire speech sample. This particularly yields an extensive context especially for frame-level acoustic feature sequences, leading to a considerable increase in computational complexity. Numerous efforts have been made to address this challenge, with one notable approach being the utilization of a state space model (SSM). SSM-based approaches [10, 11, 12, 13] have been developed to handle sequential data across diverse tasks and modalities. By integrating a time-varying mechanism into SSMs, a new model named Mamba [14] has been proposed and shown outstanding performance in CV [15, 16] and NLP [17].
@@
However, in the field of speech processing, despite some attempts to replace transformers with Mamba [18, 19, 20], the results have not been as satisfactory as expected. In [18], each Mamba is directly employed as a substitute for a Transformer within a dual-path framework for speech separation. [19] proposed SPMamba for speech separation, where Mamba is used in conjunction with MHSA. Although these approaches achieve high performance by employing a dual-path strategy or combining Mamba with attention to form a new module, these methods negate the low time complexity of Mamba. In the domain of multi-channel speech enhancement, Mamba was implemented to enhance a SpatialNet from offline to online [20] yet underperformed the vanilla version. Since different speech tasks focus on various characteristics of a speech signal (e.g., speaker, language, emotion), they generally require different levels of information. However, existing approaches have mostly investigated speech enhancement and separation tasks, which focus primarily on the low-level information within a speech signal. Therefore, it is still unclear how to efficiently employ Mamba for other speech tasks, such as speech recognition and spoken language understanding, which require high-level semantic information within the speech signal.
@@
In this paper, we provide solutions for applying Mamba to different speech tasks based on their varying information requirements (in different abstraction levels [21]) using speech recognition and speech enhancement as examples. Our research first proposes and compares two bidirectional Mamba (BiMamba) structures, being external BiMamba (ExtBiMamab) and inner BiMamba (InnBiMamba). Experiments suggest that a bidirectional design can enhance the capability of Mamba to model global dependencies within the features of a speech signal. Mamba and BiMamba models are then evaluated independently or as replacements for MHSA in Transformer and Conformer models across multiple datasets. We demonstrate that the proposed BiMamba modules require additional nonlinearity to learn high-level semantic information in speech tasks, serving as an alternative to MHSA is thus an optimal approach to apply Mamba in this scenario.
2 Preliminary
The State Space Model (SSM) based approaches, including the Structured State Space Sequence Model (S4) [11] and Mamba [14], are derived from continuous systems. S4 facilitates the transformation of a one-dimensional function or sequence, to , through an intermediate state in . The process leverages from as the evolution matrix and from , from as the input and output mapping matrices, respectively.
The fundamental equations are represented as:
In their discrete forms, parameter D can be represented as the residual connection in Neural Network, S4 and Mamba introduce a scaling parameter , transforming the continuous matrices into the discrete matrices respectively. This transformation commonly utilizes the Zero-Order Hold (ZOH) method, defined by:
Consequently, with the discretization into , the equation (1) is adjusted for a discrete timestep , given as:
The model then computes the final output from convolution via
The matrix, as described above, is commonly initialized with a HiPPO matrix [22] or a Diagonal matrix [12] to adeptly capture long-range dependencies. Mamba enhances S4 by integrating a time-varying mechanism, which enlarges the dimensions of matrices and to , and modifies and to .
@@ 2 Preliminary
Although these modifications have enhanced model performance, granting it "selective" capabilities, they do not alter the inherent nature of SSMs, which operate in a unidirectional manner such as RNNs [23, 24]. This is not an issue in training large language models, as many of these models are trained in an autoregressive manner [25, 26]. However, for non-autoregressive speech models, we require a module with causal capabilities similar to attention. Thus, finding a suitable method to address this issue is essential. Moreover, the observation that Mamba’s matrix randomization and real-valued diagonal initialization perform equivalently suggests that Mamba’s ability to delineate dependencies between inputs, compared to S4, needs further enhancement.
3 Investigating Mamba in Speech Processing
3.1 Bidirecional processing
@@
The original Mamba performs causal computations in a unidirectional manner, using only historical information. However, in speech tasks, the model is provided with the complete speech signal. Therefore, Mamba requires bidirectional computations, as employed in the MHSA module, to capture global dependencies within the features of the input signal. In this paper, we explored two bidirectional strategies for Mamba in speech tasks, i.e., inner bidirectional Mamba (InnBiMamba) [15] and external bidirectional Mamba (ExtBiMamba) as shown in Figure 2.
@@
Inner Bidirectional Layer (InnBiMamba). We first explore the inner bidirectional Mamba (InnBiMamba) [15] for speech tasks as detailed in Figure 1(a). Here, two SSM modules share the same input and output projection layers. The process feeds the input forward into one SSM module, while reversing the input along the time dimension before feeding it into the other SSM module. The output of the backward SSM module is reversed back before being combined with the output of the forward SSM module. The combined output then passes through the output projection layer.
External Bidirectional Layer (ExtBiMamba). In addition, we also propose a simpler and more straight bidirectional modeling strategy, i.e., external bidirectional Mamba (ExtBiMamba). Unlike the InnBiMamba layer, the ExtBiMamba layer involves different input and output projection layers in forward and backward Mamba layers, respectively, as illustrated in Figure 1(b). The input is fed into the forward Mamba layer and the backward Mamba layer processes the reversed input. The outputs of two Mamba layers are fused with addition operation and a residual connection is applied around the ExtBiMamaba layer. Please refer to Appendix A for detailed descriptions of two algorithms.
3.2 Task-aware model designs
@@
To capture information of various abstraction levels, we progressively explored three structures for increasing nonlinearity capability. As depicted in Figure 2(a), the first strategy uses the Mamba/BiMamba layers independently (i.e., as a direct replacement for the transformer layer) to construct the Mamba/BiMamba model. The second approach employs the Mamba/BiMamba layer to replace the MHSA modules within the Transformer, where the feed-forward net (FFN) and layernorm modules are used to provide nonlinearity. The third replaces the MHSA modules with Mamba/BiMamba layers in the Conformer, which is a variant of the Transformer employing a convolutional layer after each MHSA designed to additionally capture local information.
4 Experimental Setup
4.1 Speech Enhancement
@@
Datasets. We follow the studies [28, 29] and employ the clean speech clips from LibriSpeech train-clean-100 corpus [30] for training, comprising clips spoken by 251 speakers. The noise recordings are collected from the following datasets [31], i.e., the noise data of the MUSAN datasets [32], the RSG-10 dataset [33] (voice babble, F16, and factory welding are excluded for testing), the Environmental Noise dataset [34, 35], the colored noise set (with an value ranging from -2 to 2 in increments of 0.25) [36], the UrbanSound dataset [37] (street music recording no 26 270 is excluded for testing), the QUT-NOISE dataset [38], and the Nonspeech dataset [39]. Please refer to Appendix C.1 for the detailed experiment setup.

### v6

@@ Abstract
Transformer and its derivatives have achieved success in diverse tasks across computer vision, natural language processing, and speech processing. To reduce the complexity of computations within the multi-head self-attention mechanism in Transformer, Selective State Space Models (i.e., Mamba) were proposed as an alternative. Mamba exhibited its effectiveness in natural language processing and computer vision tasks, but its superiority has rarely been investigated in speech signal processing. This paper explores solutions for applying Mamba to speech processing by discussing two typical speech processing tasks: speech recognition, which requires semantic and sequential information, and speech enhancement, which focuses primarily on sequential patterns. The experimental results confirm that bidirectional Mamba (BiMamba) consistently outperforms vanilla Mamba, highlighting the advantages of a bidirectional design for speech processing. Moreover, experiments demonstrate the effectiveness of BiMamba as an alternative to the self-attention module in the Transformer model and its derivates, particularly for the semantic-aware task. The crucial technologies for transferring Mamba to speech are then summarized in ablation studies and the discussion section, offering insights for extending this research to a broader scope of tasks.
Index Terms:
Speech Recognition, Speech Enhancement, Mamba, Selective State Space Model, Self-AttentionI Introduction
@@
Transformer-based models [1] have shone brightly across various domains in machine learning, including computer vision (CV) [2, 3, 4], natural language processing (NLP) [5, 6, 7], and speech processing [8, 9, 10, 11]. This success is linked to the multi-head self-attention (MHSA) module, which facilitates the representation of intricate data structures within a specific context window. However, the self-attention mechanism encounters a challenge with computational complexity, which grows quadratically as the size of the context window increases. In speech tasks, the window typically encompasses an entire speech signal, leading to substantial computational demands, particularly for frame-level acoustic feature sequences.
@@
To address this challenge, state space models (SSMs) have emerged as a promising alternative [12, 13, 14, 15]. By integrating a time-varying mechanism into SSMs, a selective SSM named Mamba [16] has been proposed and shown outstanding performance to be effective in CV [17, 18] and NLP [19]. However, in the field of speech processing, despite some attempts to replace Transformers with Mamba [20, 21, 22], the results have not been as satisfactory as expected. In [20], each Mamba is directly employed as a substitute for a Transformer within a dual-path framework for speech separation. [21] proposed SPMamba for speech separation, where Mamba is used in conjunction with MHSA. Although these approaches achieve high performance by employing a dual-path strategy or combining Mamba with attention to form a new module, these methods negate the low time complexity of Mamba. In the domain of multi-channel speech enhancement, Mamba was implemented to enhance a SpatialNet from offline to online [22] yet underperformed the vanilla version.
Since different speech tasks focus on various characteristics of a speech signal (e.g., speaker, language, emotion), they generally require different levels of information. However, existing approaches have mostly investigated speech enhancement and separation tasks, which focus primarily on the low-level information within a speech signal. Therefore, it is still unclear how to efficiently employ Mamba for other speech tasks, such as speech recognition and spoken language understanding, which require high-level semantic information within the speech signals.
@@
In this paper, we explore solutions for applying Mamba to different speech tasks based on their varying information requirements (in different abstraction levels [23]) using speech recognition and speech enhancement as examples. We first introduce and compare two bidirectional Mamba (BiMamba) structures, external BiMamba (ExtBiMamba) and inner BiMamba (InnBiMamba), in speech tasks. The experiments confirm that a bidirectional design can enhance the capability of Mamba to model global dependencies within the features of a speech signal. Mamba and BiMamba models are then evaluated independently or as replacements for MHSA in Transformer and Conformer models. We demonstrate that the proposed BiMamba modules require additional nonlinearity to effectively learn high-level semantic information in speech tasks. Therefore, using BiMamba as an alternative to MHSA presents an optimal approach for applying Mamba in this scenario.
II Preliminary
II-A State Space Models and Mamba
The State Space Model (SSM) based architectures, such as the Structured State Space Sequence (S4) Model [13] and Mamba [16], are typically inspired by the continuous linear time-invariant (LTI) systems. It maps a sequence to by leveraging a hidden state . Formally, the mapping process of SSMs can be formulated as follows:
where , , , and (optional) represent the continuous SSM parameters [16].
SSM Discretization. In Mamba and S4, a discretization process is applied to continuous-time SSMs for the integration into practical deep neural architectures. A timescale parameter is introduced to transform the continuous matrices to their discrete counterparts . The zero-order hold (ZOH) is a commonly used approach for the transformation, which is formulated as follows:
The approximation of refined using the first-order Taylor series is given by . Thus, the discrete SSM is expressed as:
The output sequence can be computed simultaneously through a global convolution operation as follows:
where is a convolution kernel derived from the SSM, is the length of the input sequence, and denotes the convolution operation. Since the parameters , , and remain constant with respect to input contents (temporal dynamics), previous S4 models struggle to effectively capture contextual information.
Selective SSM. Mamba [16] improves S4 by incorporating a selective mechanism that enables adaptive parameter adjustment based on input characteristics, presenting the selective state space model. The parameters , , and are computed as functions of the input. Given the input sequence and the output sequence , the selective SSM handles each channel independently through the hidden state , and the formulation can be written as follows [24]:
where , , and are derived from the input, and denotes the Hadamard product. Mamba generates the parameters , , and using the following projections: , , where , , and are projection matrices, and refers to . The reduction ratio of the two-layer linear projection for is set to 16 by default [16]. Equation 5 represents the exact selective SSM used in Mamba, with only slight formatting adaptations.
The selective mechanism features the dynamic parameters, allowing the model to learn context-aware representations and effectively filter out irrelevant information. However, it poses a computational challenge: the convolution kernels become input-dependent, preventing parallel convolution operations. To overcome this problem, Mamba employs a hard-aware algorithm to speed up computation, leveraging three classical techniques, i.e., parallel associative scan (also called parallel prefix-sum) [25], kernel fusion, and recomputation. Researchers observed that each state represents the whole compression of the previous states, enabling the direct computation of new states from previous ones. This suggests that the execution order of operations is independent of their associated attributes. Thus, Mamba employs the selective scan algorithm by computing sequences in segments, iteratively combining them, and integrating parameters that are conditioned on the input. In addition, GPUs consist of numerous processors capable of highly parallel computations. Mamba leverages the GPU’s high-bandwidth memory (HBM) and fast SRAM I/O to avoid frequent SRAM writes to HBM through kernel fusion. Concretely, the discretization and recursive operations are performed in the higher-speed SRAM memory, and the output is then written back to the HBM. During backpropagation, the intermediate states are not stored but are recomputed when inputs are loaded from the HBM to the SRAM.
@@ II-A State Space Models and Mamba
Although these modifications have enhanced model performance, granting it “selective” capabilities, they do not alter the inherent nature of SSMs, which operate in a unidirectional manner such as RNNs [26, 27]. This is not an issue in training large language models, as many of these models are trained in an autoregressive manner [28, 29]. However, for non-autoregressive speech models, we require a module with non-causal capabilities similar to attention. Thus, finding a suitable method to address this issue is essential. Moreover, the observation that Mamba’s matrix randomization and real-valued diagonal initialization perform equivalently suggests that Mamba’s ability to delineate dependencies between inputs, compared to S4, needs further enhancement.
II-B Mamba for Audio Processing
Recent works have explored the use of SSMs and Mamba methods in speech enhancement and separation. Distilled Audio SSM (DASS) applied knowledge distillation with SSMs and outperformed transformer-based models with processing sequences up to 2.5 hours long [30]. Audio Mamba (AuM) [31] and its variant for audio tagging [32] achieved comparable or better performance than Audio Spectrogram Transformers (AST) with roughly one-third of the parameters. Meanwhile, Self-Supervised Audio Mamba (SSAM) [33] established that unidirectional Mamba blocks were particularly effective for masked spectrogram modeling, consistently outperforming transformer-based approaches while better handling varying sequence lengths. [34] proposed a SEMamba method for speech enhancement, which improved state-of-the-art performance by integrating time-frequency Mamba into advanced SE architecture. These advances collectively demonstrate the potential of Mamba architectures as an efficient and scalable alternative to transformers for audio processing tasks.
[35] conducted comprehensive experiments to compare Mamba encoder and decoder with existing methods across various tasks, including TTS, ASR, SLU, and SUMM. Mamba encoder and decoder are introduced. [36] incorporates the Mamba module in state-of-the-art methods of speech separation, ASR, and TTS, comparing them with the vanilla models in their respective tasks. [34] proposed a SEMamba method for speech enhancement, which improved state-of-the-art performance by integrating time-frequency Mamba into advanced SE architecture. Our work, similar to [35] and [36], explores the use of Mamba in speech processing tasks. However, we specifically investigate Mamba’s ability to learn speech information at various levels of abstraction, positioning it as an alternative to the self-attention module within Transformer and its derivatives. Beyond experiment-based analysis, we delve into the reasons behind the independent Mamba model’s lower performance in ASR by examining the impact of nonlinearity.
III Investigating Mamba in Speech Processing
III-A Bidirectional processing
@@
The original Mamba performs causal computations in a unidirectional manner, using only historical information. However, in speech tasks, the model is provided with the complete speech signal. Therefore, Mamba requires bidirectional computations, as employed in the MHSA module, to capture global dependencies within the features of the input signal. In this paper, we explored two bidirectional strategies for Mamba in speech tasks, i.e., inner bidirectional Mamba (InnBiMamba) from vision Mamba [17] and external bidirectional Mamba (ExtBiMamba) as shown in Figure 1.
@@
InnBiMamba. We first explore the inner bidirectional Mamba (InnBiMamba) from Vision Mamba [17] for speech tasks as detailed in Figure 1(a). Here, two SSM modules share the same input and output projection layers. The process feeds the input forward into one SSM module, while reversing the input along the time dimension before feeding it into the other SSM module. The output of the backward SSM module is reversed back before being combined with the output of the forward SSM module. The combined output then passes through the output projection layer.
ExtBiMamba. We next propose a simpler and more straight bidirectional modeling strategy, i.e., external bidirectional Mamba (ExtBiMamba). The ExtBiMamba design is motivated by several key considerations in bidirectional sequence modeling. While existing approaches often rely on internal bidirectional mechanisms, our external approach offers distinct advantages. First, by maintaining separate input and output projections for forward and backward directions, ExtBiMamba allows each direction to learn direction-specific feature transformations that are optimal for processing the sequence in that particular order. This is particularly valuable for modeling contextual dependencies in speech signals, where certain acoustic patterns may be more interpretable when traversed in one direction rather than the other.
The separation of forward and backward paths also enables more explicit modeling of temporal dependencies. The forward pathway captures the natural progression of acoustic events, while the backward pathway can effectively utilize future information when analyzing the sequence in reverse. This is critical for tasks that rely on contextual and sequential information, like ASR, speech enhancement, or emotion recognition. Compared to the InnBiMamba, the proposed ExtBiMamba design allows the forward and backward directions to learn complementary and non-redundant feature spaces. This separation enhances the capacity of BiMamba to represent complex, hierarchical structures in speech and contributes to improved performance in tasks requiring fine-grained temporal understanding. Detailed algorithms of these two methods are included in Section III-C.
III-B Task-aware model designs
@@
To capture information of various abstraction levels, we progressively explored three structures for increasing nonlinearity capability. As depicted in Figure 2(a), the first strategy uses the Mamba/BiMamba layers independently (i.e., as a direct replacement for the transformer layer) to construct the Mamba/BiMamba model. The second approach employs the Mamba/BiMamba layer to replace the MHSA modules within the Transformer, where the feed-forward network (FFN) and layer normalization are used to provide nonlinearity. The third replaces the MHSA modules with Mamba/BiMamba layers in the Conformer, which is a variant of the Transformer employing a convolutional layer after each MHSA designed to additionally capture local information.
III-C Algorithms of Bidirectional Mamba modules
Algorithms 1 and 2 illustrate the workflows of InnBiMamba and ExtBiMamba, respectively. The main differences are highlighted in violet. For the InnBiMamba layer, the linear input projections ( and ) and the linear output projection are shared across forward and backward operations. In contrast, the ExtBiMamba layer uses different input linear projections ( and ) and output linear projections () across forward and backward operations. The dimensionality of the matrices is defined as follows: represents the training batch size, denotes the sequence length, denotes the dimension of the input hidden state, and denotes the expanded hidden state dimension, and denotes the SSM state dimension.
IV Experimental Setup
IV-A Speech Enhancement
@@
Datasets. We follow previous works [37, 38] and employ the clean speech data from the LibriSpeech train-clean-100 corpus [39] as the training set, containing speech clips spoken by speakers. The noise recordings are collected from the following datasets [40], i.e., the noise data of the MUSAN datasets [41], the RSG-10 dataset [42] (voice babble, F16, and factory welding are excluded for testing), the Environmental Noise dataset [43, 44], the colored noise set (with an value ranging from -2 to 2 in increments of 0.25) [45], the UrbanSound dataset [46] (street music recording no 26 270 is excluded for testing), the QUT-NOISE dataset [47], and the Nonspeech dataset [48].
Noise recordings that exceeds seconds in duration are split into clips of seconds or less. This yields noise clips, with each clip less than or equal to seconds in duration. For validation experiments, clean speech and noise clips (without replacement) were randomly drawn from the aforementioned clean speech and noise sets and mixed to generate a validation set of noisy clips, where each clean speech clip was degraded by a random section of one noise clip at a random SNR level (sampled between and dB, in 1 dB steps). For evaluation experiments, we employed four real-world noise sources (excluded from the training set) including two non-stationary and two colored ones. The two non-stationary noise sources were the voice babble from the RSG-10 noise dataset [42] and street music from the Urban Sound dataset [46]. The two colored noise sources were F16 and factory welding from RSG-10 noise dataset [42]. For each of the four noises, we randomly picked twenty clean speech clips (without replacement) from the test-clean-100 of LibriSpeech corpus [39] and degraded each clip with a random section of the noise clip at the five SNR levels, i.e., . This generated noisy mixtures for evaluation.
Feature Extraction. All audio signals are sampled at a rate of 16 kHz. We employ a 512-sample (32 ms) long square-root-Hann window with a hop length of 256 samples (16 ms), to extract a 257-point single-sided STFT spectral magnitude as the input to the neural models [40].